X is continuous uniform (1,7) while Y is exponential with mean 2. If the variance of (X+2Y) is 20, find the correlation coefficient of X and Y
Answer #1

Given

Var(X+2Y) = 20

X~ Unif(1,7)

Y~ Exp(\lambda)

E(Y) =\frac{1}{\lambda }=2

\Rightarrow \lambda =\frac{1}{2}

We first compute the variance of X and Y.

For Unifrom distribution,

Var(X)=\frac{\left ( b-a \right )^{2}}{12}

                   =\frac{\left ( 7-1 \right )^{2}}{12}=3

and for exponential distribution,

Var(Y)=\frac{1}{\lambda ^{2}}=\frac{1}{ \left ( \frac{1}{2} \right )^{2}}

                   =4

Now,

Var(X+2Y)=Var(X)+4Var(Y)+4Cov(X,Y)

\Rightarrow Var(X)+4Var(Y)+4Cov(X,Y)=20

\Rightarrow 3+4\times4+4Cov(X,Y)=20

\Rightarrow 4Cov(X,Y)=20-19

\Rightarrow 4Cov(X,Y)=1

\Rightarrow Cov(X,Y)=\frac{1}{4}

Answer: 1/4